# Entangled photon pair excitation and time-frequency filtered multidimensional photon correlation spectroscopy as a probe for dissipative exciton kinetics

### v1

@@ Abstract
In molecular aggregates, multiple delocalized exciton states interact with phonons, making the state-resolved spectroscopic monitoring of dynamics challenging. We propose a protocol that combines photon-entanglement-enhanced narrowband excitation of two-exciton states with time-frequency-filtered two-photon coincidence counting. It can alleviate bottlenecks associated with probing exciton dynamics spread across multiple spectral and temporal windows. We demonstrate that non-classical correlations of entangled photon pairs can be used to prepare narrowband two-exciton population distributions, circumventing transport in mediating states. The distributions thus created can be monitored using time-frequency-filtered photon coincidence counting, and the pathways contributing to photon emission events can be classified by tuning filtering parameters. Numerical simulations for a light-harvesting aggregate highlight the ability of this protocol to achieve selectivity by suppressing or amplifying specific pathways. Combining entangled photonic sources and multidimensional photon counting allow promising applications to spectroscopy and sensing.
@@ I Introduction
In photosynthetic systems, light-harvesting complexes are critical components that facilitate efficient photon energy conversion through exciton formation and transport. Investigating the structural and functional mechanisms of exciton dynamics provides essential insights for engineering biomimetic energy-harvesting systems. In these systems, aggregated chromophores serve as the energy harvesters. The aggregation of individual chromophores results in delocalized electronic excitations: molecular excitons which are primarily of the Frenkel type. The excitons, often generated through photo-excitation, participate in multiscale relaxation and dephasing kinetics, due to the interaction with phonons [16, 1, 30], ultimately leading to the primary charge separation processes. Previous investigations have sought to monitor the role of coherent energy transfer that results from transition dipole-dipole interactions. The later can be probed via multi-pulse four-wave mixing techniques, popularly known as two-dimensional spectroscopy. The investigation multi-exciton states, where signatures of exciton-exciton correlations can be found, have comparatively received less attention. The early attributes of such correlations can already be seen in the two-exciton state energetics. Prior investigations have demonstrated that various modalities of two-photon absorption spectroscopy, resonant and off-resonant, can spectroscopically monitor these resonances [50, 51, 20, 15, 35, 54, 49, 7, 53, 26, 25, 33, 48, 21, 45, 29, 4]. Probing state-resolved dynamics in the two-exciton manifolds is complicated by the high density of states that take part in phonon-induced dissipation spanning multiple time and energy scales. Over the past several years, a series of proposals, limited to atomic, and simple molecular systems have underscored the advantages of employing entangled photon sources to improve accessibility of doubly excited states while scaling favorably with the probe intensity [46, 23, 44, 40, 14, 22, 8, 18, 19, 17, 43, 52, 32, 5, 34]. These proposals were complemented by examinations of the role of photon entanglement in the context of the reported enhancements [42, 28, 27, 39, 41, 24, 3, 6]. Parallel theoretical investigations have established that exploiting the non-classical time-frequency correlations of photonic states may yield superior signal resolution in spectroscopic measurements compared to classical counterparts [10, 8].
To monitor the fate of two-exciton states, multiphoton fluorescence intensity correlation measurements have been proposed [34, 31, 38, 11, 13]. In a series of work [11, 13], the concept of multidimensional signal measurements was extended to include coincidence counting measurements. This theoretical framework promises to transparently link information derived from photon correlation measurements to molecular events. The time and frequency filtered $n$-photon coincidence counting measurements provide a $2n$-dimensional parameter space for the events to be classified. However to firmly establish their role as a routine spectroscopic tool in extended molecular systems, it requires a protocol combining them with efficient initial state preparation strategies.
The two-exciton states in molecular aggregates, such as light-harvesting complexes, are typically prepared via the manifold of one-exciton states, which are subject to phonon-induced transport and dephasing. Furthermore, as the number of chromophoric sites increases, the number of one-exciton and two-exciton states increases as well. As a consequence, the number of interfering optical transitions may increase dramatically. Additionally, phonon-induced transport redistributes the population among the participating states during the dynamics.
In this work, we demonstrate that state-selective entangled photon excitation and time-frequency-filtered multidimensional correlation spectroscopy can serve as an efficient probe for dissipative exciton kinetics. Specifically, this approach helps unravel two-exciton transport kinetics by monitoring cascaded photon emission events, tracking the correlation between these events, and benchmarking the nature of the exciton wavepacket created by entangled photon excitation.
We focus on the light-harvesting complex LHCII, the most abundant antenna complex in plants and an important constituent of Photosystem II. It exists as a trimeric structure composed of structurally rigid monomeric aggregates; each monomer contains of several chlorophylls whose lowest energy levels are spectroscopically accessible. Exciton dynamics in LHCII exhibit significant complexity, particularly regarding the behavior of higher-order excitations. The combined protocol is shown to provide mechanistic insight into the cooperative exciton kinetics of LHCII, even in the presence of spectral weight redistribution caused by mode mixing and relaxation.
Figure 1: (A) Illustration of the proposed protocol, consisting of the driving, evolution, and detection stages. The process is assisted by entangled photon pairs and temporal-spectral filtering of emitted photons. In (B), three plausible snapshots of the dynamics (I)-(III), featuring increasingly complex pathways, that may be monitored using the protocol is displayed. In the simulation, the exciton transport within both manifolds is accounted for. Since these pathways may be simultaneously present in the dynamics, the protocol allows for selective probing and aids in constructing a mechanistic narrative.
@@
In the following, in section II, we introduce the exciton-phonon Hamiltonian of the light-harvesting complex LHCII, which serves as our test system, and the photonic sources. This is complemented by the Appendix A. In the subsequent three subsections of Sec. III (III.1III.3), we detail the theoretical framework and numerical simulations of the combined protocol. The manuscript concludes in Sec. IV with a summary of our findings, a discussion of future outlook, and a brief examination of the current limitations.
@@ II Driven dissipative one- and two-exciton kinetics
The dissipative one- and two-exciton dynamics of the light-harvesting complex LHCII are simulated using the Frenkel exciton model within the Heitler-London approximation. The model is constructed by considering the lowest three electronic energy states of each of the chlorophyll molecules. The Hamiltonian (with $\hbar=1$) is given by,
@@
Here, the on-site excitation energy, $E_{m}$, and inter-site hopping, $J_{mn}$, are accounted for by the first term in the Hamiltonian. The Coulomb interaction, responsible for the inter-site hopping, results in the delocalization of the exciton states across the sites. The higher-order excitonic nonlinearities that shift exciton energy states are given by $U_{m}^{(1)}$ and $U_{mn}^{(2)}$. The exciton creation (annihilation) operators, $B_{m}^{\dagger}$ ($B_{m}$), follow the commutation relation $B_{n},B_{m}^{\dagger}]=\delta_{mn}(1-\eta(B_{m}^{\dagger}B_{m})^{2})$, where $\eta=3/2$, while the phonon creation (annihilation) operators, $b_{j}^{\dagger}$ ($b_{j}$), follow $b_{i},b_{j}^{\dagger}]=\delta_{ij}$. The two-exciton Hamiltonian can be constructed as $H_{mnkl}^{(2)}=H_{mn}^{(1)}\delta_{kl}+\delta_{mn}H_{kl}^{(1)}$, where $H_{mn}^{(1)}$ is given by the first line of the Eq. II. For the LHCII, the number of chromophoric sites is $N_{s}=14$, which yields $N_{g}=1$ ground state, $N_{e}=14$ one-exciton states, and $N_{f}=105$ two-exciton states, composed of $14$ overtone and $91$ combination states. The excitonic eigenstates are obtained by the exact diagonalization of the exciton-number conserving Hamiltonian blocks as:
@@
These transformation coefficients, $T^{(n)}$ (for the nthn-th manifold) define the spectral weights associated with the delocalized eigenstates. It is noteworthy that the energies of the two-exciton overtone and combination states, dependent on the nonlinearity terms $U_{m}^{(1)}$ and $U_{mn}^{(2)}$ respectively, deviate from the sum of two one-exciton states. As a consequence, the relative ordering of states within the two-exciton manifold is contingent upon two competing sets of parameters, governing exciton delocalization and localization. The latter influence, via excited state absorption processes, all three pathways responsible for the two-photon absorption events addressed in this article. The Green’s functions (valid for both one and two exciton manifold) describing can be represented in a compact form as follows
@@
The optical properties of exciton aggregates consisting of multiple three level emitters are primarily dictated by the nature of intermolecular interactions and two-exciton nonlinearities [26, 2, 1]. The interaction with photonic sources, within the rotating wave approximation, can be described by the following Hamiltonian
$H_{\mathrm{int}}(t)=EV^{\dagger}+E^{\dagger}V=\sum_{j,pq\in\{ge,ef\}}\sqrt{\frac{2\pi\omega_{j}}{V}}$
@@
where $B_{pq}(t)=d_{pq}|p\rangle\langle q|(\text{with properties}\,\,\{p,q\}\in g,e,f$, and $a_{j}$ are the exciton and photon annihilation operators, describing inter-manifold exciton transition and excitations in the photonic modes, respectively.
Figure 2: (A) Diagrammatic illustration (using the Schwinger-Keldysh loop representation) of the pathways involved in the proposed protocol, consisting of the driving, evolution, and emission stages. (B)-(C) Diagrammatic illustration (using the Liouville superoperator representation, suitable for dissipative systems) of the pathways involved in filtered emission (B) and excitation (C). The combinatorial pairing of five possible excitation diagrams and four possible emission diagrams gives rise to all possible pathways that contribute to the final signal. The corresponding key equations are presented in Section III.
For our case, the deployed photonic sources are entangled photon pairs generated via spontaneous parametric down-conversion (SPDC) in the weak-down-conversion limit. In a typical scenario, a birefringent crystal is pumped with a classical pump that precedes to the generation of correlated photon pairs with frequencies $\omega_{s},\omega_{i}$. It involves two field-matter interactions with the incoming classical field modes and four interactions with the outgoing signal-idler fields. The correlation properties of the emitted photon field is determined by the physical properties of the SPDC crystal [44, 47, 23, 46, 12, 8, 9]. The incoming entangled photon states can be described by the two photon state of the photon field
@@
The entangled photon field correlation function is obtained as,
$\langle E^{{\dagger}}(\omega_{b}^{\prime})E^{{\dagger}}(\omega_{a}^{\prime})E(\omega_{b})E(\omega_{b})\rangle=F^{*}(\omega_{b}^{\prime},\omega_{a}^{\prime})F(\omega_{a},\omega_{b})$ (6)
where,
$F(\omega_{a},\omega_{b})$ $=\alpha A_{p}(\omega_{p})\Big\{\mathrm{sinc}[\phi(\omega_{a},\omega_{b})]$
$\exp{(i\phi(\omega_{a},\omega_{b}))}+\omega_{1}\leftrightarrow\omega_{2}\Big\}\text{with}$
$\phi(\omega_{a},\omega_{b})$ $=(\omega_{a}-\omega_{1})\tilde{T}_{1}/2+(\omega_{b}-\omega_{1})\tilde{T}_{2}/2$ (7)
where $\alpha$ governs the down-conversion efficiency. The temporal entanglement parameter of the photon pair sources is given by $\tilde{T}_{\mathrm{ent}}=\tilde{T}_{2}-\tilde{T}_{1}$ establish the upper bound for the interaction time interval. The transform-limited field profile of the SPDC pump, $A_{p}(\omega_{a},\omega_{b})$ is taken as a Gaussian, given by, $A_{p}(\omega_{a},\omega_{b})=E_{0}\sqrt{\pi/\Gamma_{p,0}}\exp{[-(\omega-\omega_{a}-\omega_{b})^{2}/4\Gamma_{p,0}]}$, where $\Gamma_{p,0}$ is the temporal width. The quantity is dependent on the time delays acquired by the signal and idler photons during the propagation within the nonlinear crystal. In contrast to entangled photon pairs, the correlation function of coherent laser fields would have factorized into the product of four amplitude functions, i.e.
$\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
$=A^{*}_{4}(\omega_{4})A^{*}_{3}(\omega_{3})A_{1}(\omega_{2})A_{1}(\omega_{1})$ (8)
and the signal would have scaled with the square of the incoming field intensity.

### v2

@@ Abstract
In molecular aggregates, multiple delocalized exciton states interact with phonons, making the state-resolved spectroscopic monitoring of dynamics challenging. We propose a protocol that combines photon-entanglement-enhanced narrowband excitation of two-exciton states with time-frequency-filtered two-photon coincidence counting. This approach alleviates bottlenecks associated with probing exciton dynamics spread across multiple spectral and temporal windows. We demonstrate that non-classical correlations of entangled photon pairs can be used to prepare narrowband two-exciton population distributions, thereby circumventing transport in mediating one-exciton states. The evolution of these population distributions and cascading transitions can be monitored using time-frequency-filtered photon coincidence counting. Numerical simulations for a light-harvesting aggregate highlight the ability of this protocol to achieve selectivity by suppressing or amplifying specific pathways. Combining entangled photonic sources with multidimensional photon counting allows promising applications in spectroscopy and sensing.
@@ I Introduction
In interacting molecular aggregates, optical probing of dissipative exciton kinetics and the extraction of state-resolved spectroscopic information are inherently difficult. The high density of exciton states and exciton-phonon interactions lead to dissipative dynamics that span multiple time and energy scales. Molecular aggregates are found in naturally occurring photosynthetic systems. Within these photosynthetic systems, light-harvesting complexes are critical components that facilitate efficient photon energy conversion through exciton formation and transport. Investigating the structural and functional mechanisms of exciton dynamics provides essential insights for engineering biomimetic light-harvesting systems. [19, 1, 32]. In these systems, the aggregation of individual chromophores results in delocalized electronic excitations. These molecular excitons are primarily of the Frenkel type. The generated excitons participate in relaxation and dephasing, leading to primary charge-separation processes. Signature of exciton-exciton interactions and correlations, which are typically hidden in the dynamics of multi-exciton states were investigated [38, 35, 53]. The early attributes of such correlations can already be seen in the energy distributions of two-exciton states, which deviate from the sum of two one-exciton energies. Prior studies have demonstrated that various modalities of two-photon absorption spectroscopy can monitor the associated resonances. [54, 55, 23, 57, 29, 58, 8].
In recent years, several proposals restricted to atomic and simple molecular systems have underscored the advantages of employing entangled photon pairs to probe two-exciton states [51, 25, 49, 44, 18, 24, 9, 21, 22, 20, 48, 56, 34, 4, 6]. These proposals were complemented by critical examinations of the role of photon entanglement in enhancing two-photon absorption [47, 31, 30, 43, 46, 26, 2, 7]. Theoretical investigations have also explored the role of non-classical time-frequency correlations in obtaining superior resolution in condensed-phase spectroscopies. [14, 9].
Alongside, previous studies have proposed novel two-photon fluorescence intensity correlation measurements that can monitor the properties of two-exciton states [36, 33, 42, 15, 17, 50]. In a series of theoretical works, time-frequency filtered photon coincidence counting was proposed as a technique that may provide spectroscopic information similar to that of two-dimensional spectroscopy, but at different timescales [15, 17].
The optical excitation of two-exciton states in molecular aggregates, such as light-harvesting complexes, typically occurs via a manifold of one-exciton states subjected to population transport and dephasing. As a result, the population is redistributed among several states which may not be have highest propensity of transitions to the target two-exciton states. As the number of chromophoric sites increases, the number of one-exciton and two-exciton states increases as well. Consequently, the number of interfering optical transitions increase dramatically. The final two-exciton population also undergoes transport, dispersing any initially created narrowband two-exciton distribution. In this work, we demonstrate that two-photon excitation using entangled photon pairs may achieve the creation of narrow-band population distribution. We also show that the time-frequency-filtered two-photon counting can characterize cascaded photon emission events, resulting two-dimensional correlation plots. It constitutes a spectroscopic technique that can serve as an efficient probe for dissipative exciton dynamics.
We focus on the light-harvesting complex LHCII, the most abundant antenna complex in plants and an important constituent of Photosystem II.
Figure 1: (A) Illustration of the proposed protocol, classified as a three-stage process: preparation of the two-exciton population, dissipative evolution of the two-exciton population, and time-frequency filtered detection of cascaded photon emission. The action of the time filter (governed by $F_{\bar{t}}^{(j)}$) precedes the action of the frequency filter (governed by $F_{\bar{\omega}}^{(j)}$) in both cases. The vertical straight arrows denote interaction with the driving photon modes, and the wiggly arrows denote interaction with the photon modes of the detector. In (B), three exemplery dynamical pathways are shown. The tilted double-ended arrows denote intra-manifold exciton transport. In (I), both two- and one-exciton transport preserves the narrowband nature of the created population distribution. In (II), during the excitation stage, the participation of one-exciton transport alters the target two-exciton population distribution. The population distribution disperses due to two-exciton transport, which results in broadband cascaded emissions. In (III), the narrowband two-exciton population distribution funnels to an even narrower distribution, but the subsequent one-exciton transport erases the memory. It lead leads to a narrowband emission followed by a broadband emission. The filtering functions assist in constructing a mechanistic narrative by classifying these pathways.
@@
It exists as a trimeric structure composed of structurally rigid monomeric aggregates; each monomer contains several chlorophylls whose lowest three energy levels are spectroscopically accessible. In Section II, we introduce the exciton-phonon Hamiltonian of the light-harvesting complex LHCII. We also the exciton-photon Hamiltonian. This is complemented by the Appendix A. In the subsequent sections, namely Sec. III (III.1III.3), we detail the theoretical framework for the combined protocol and numerical simulations. The manuscript concludes in Sec. IV with a summary of our findings and a brief commentary of the current limitations.
@@ II Driven dissipative one- and two-exciton kinetics
We simulate the one- and two-exciton dynamics of the light-harvesting complex LHCII using the Frenkel exciton model within the Heitler-London approximation. The model is constructed by considering the lowest three electronic energy states for each chromophoric site. The Hamiltonian (with $\hbar=1$) is given by,
@@
where the exciton creation (annihilation) operators, $B_{m}^{\dagger}$ ($B_{m}$), follow the commutation relation $B_{n},B_{m}^{\dagger}]=\delta_{mn}(1-\eta(B_{m}^{\dagger}B_{m})^{2})$ (with $\eta=3/2$) and the phonon creation (annihilation) operators, $b_{j}^{\dagger}$ ($b_{j}$), follow the commutation relation $b_{i},b_{j}^{\dagger}]=\delta_{ij}$.
In the above, $E_{m}$ and $J_{mn}$ describes the on-site excitation and inter-site hopping, respectively. The later arises due the Coulomb interaction and leads to the delocalization of the exciton states across the sites. The one-exciton Hamiltonian can be constructed by considering the first term, i.e., $h_{mn}^{(1)}=(E_{m}\delta_{mn}+J_{mn})B_{m}^{\dagger}B_{n}$. Upon diagonalization, we obtain the corresponding one-exciton eigenstates $|e_{j}\rangle$.
The two-exciton Hamiltonian can be constructed by considering $h_{mnkl}^{(2)}=h_{mn}^{(1)}\delta_{kl}+\delta_{mn}h_{kl}^{(1)}+U_{m}^{(1)}\delta_{mk}\delta_{nl}\delta_{mn}+U_{mn}^{(2)}\delta_{mk}\delta_{nl}$ where $h_{mn}^{(1)}$ is given above. We note that the matrix elements of the two-exciton Hamiltonian differs from the sum of two one-exciton Hamiltonian by the last two terms. These terms which contribute to two-exciton nonlinearities, represent the energetic shift of overtone and combination exciton states respectively. In a similar manner as the previous case, the two-exciton eigenstates $|f_{k}\rangle$ are obtained by diagonalization.
The parameters for the Frenkel exciton Hamiltonian were taken from [41, 40, 10]. For the LHCII, the number of chromophoric sites is $N_{s}=14$, which yields $N_{g}=1$ ground state, $N_{e}=14$ one-exciton states, and $N_{f}=105$ two-exciton states, composed of $14$ overtone and $91$ combination states. The eigenstates can be expressed as
@@
where $T^{(1)}_{j,m}$ and $T^{(2)}_{k,mn}$ denote the matrix elements of the diagonalization matrices mentioned above. These matrix elements determine the spectral weights associated with the eigenstates which are delocalized over the site basis.
Figure 2: One- and two-exciton energies are displayed in the left and right panels, respectively. Nonuniform energy distributions are visible in both manifolds. Simulations in Section III.1 refer to the excitation of two-exciton states $f_{07}$ ($E_{f_{07}}=30192.9\,\text{cm}^{-1}$) and $f_{83}$ ($E_{f_{83}}=31243.2\,\text{cm}^{-1}$) via $e_{07}$ ($E_{e_{07}}=15307.1\,\text{cm}^{-1}$) and $e_{09}$ ($E_{e_{09}}=15655.9\,\text{cm}^{-1}$).
The one exciton manifold spans from $E_{e_{1}}=14964.5\,\text{cm}^{-1}$ to $\omega_{e_{14}}=15934.7\,\text{cm}^{-1}$, hosting $14$ states within a spectral width of $\approx 970\,\text{cm}^{-1}$. The two-exciton manifold spans from $E_{f_{1}}=29844.3\,\text{cm}^{-1}$ to $E_{f_{105}}=31811.8\,\text{cm}^{-1}$, hosting $105$ states within a spectral width of $\approx 1968\,\text{cm}^{-1}$. The corresponding energy distributions, plotted in Fig. 2, offers a visualization of the nature of the density of states in each manifold. We note that the energetic ordering of exciton eigenstates is dependent upon two competing parameters: $J_{mn}$ that govern exciton delocalization and $U_{m}^{(1)}$, $U_{mn}^{(2)}$ that govern exciton nonlinearities.
The dissipative exciton dynamics is described by the Liouville-space Green’s function formalism. Using the exciton eigenbasis, we can express the Green’s function as follows
@@
The expression above is valid for both exciton transport and dephasing, which are captured by the first and the second terms, respectively. It is discussed further in Appendix A.
Interaction with the photonic sources, within the rotating-wave approximation, can be described by the following Hamiltonian:
$H_{\mathrm{int}}(t)=EV^{\dagger}+E^{\dagger}V=\sum_{j,pq\in\{ge,ef\}}\sqrt{\frac{2\pi\omega_{j}}{\Omega}}$
@@
In the above, we used a mode expansion of the photonic field operators, where $a_{j}$ denotes the photon annihilation operators and $\omega_{j}$ represents the mode frequencies. These mode operators follow bosonic commutation relations. The exciton annihilation operators, expressed in the eigenbasis, are denoted as $B_{pq}(t)=d_{pq}|p\rangle\langle q|(\text{with properties}\,\,\{p,q\}\in g,e,f$. The entangled photon pairs induce inter-manifold exciton transitions.
Figure 3: (A) The Schwinger-Keldysh loop diagram describing the pathways involved in the proposed protocol. The excitation and emission states involve four exciton-photon interactions each with the driving and detection photon modes. The Liouville space ladder diagrams in (B)-(C) are suitable for capturing the processes in dissipative systems. The excitation and emission pathways are depicted in (B) and (C) respectively. In the excitation stage, described in (C), diagrams (II) and (IV) depends on one-exciton transport, $|e\rangle\langle e|\rightarrow|e^{\prime}\rangle\langle e^{\prime}|$. Diagram (I) depends on inter-manifold two-exciton coherence, $|f\rangle\langle g|$, while (III) and (V) depend on intra-manifold one-exciton coherence, $|e\rangle\langle e^{\prime\prime}|$. In the emission stage, described in (B), all diagrams have the two-exciton population distribution following transport as the initial state, $|f\rangle\langle f|\rightarrow|f^{\prime}\rangle\langle f^{\prime}|$. A combinatorial pairing of five excitation and four possible emission diagrams generate all possible pathways that may contribute to the signal. Their relative contributions to the final signal is jointly determined by the four-point photon correlation functions and the filtering properties. The corresponding equations are presented in Section III.
We consider entangled photon pairs generated via spontaneous parametric down-conversion (SPDC) in the weak-down-conversion limit. In a typical scenario, a birefringent crystal is pumped with a classical laser, leading to the generation of correlated photon pairs with frequencies $\omega_{1}$ and $\omega_{2}$. The SPDC process involves two photon-matter interactions with the incoming classical field modes and four photon-matter interactions with the outgoing signal-idler fields. The correlation properties of the emitted photon field are determined by the physical properties of the SPDC crystal. [49, 52, 25, 51, 16, 9, 10].
The generated entangled two-photon states can be described as follows:
@@
The four-point correlation function of the entangled photon pairs is relevant for the creation of two-exciton states. Using the expression above, it can be evaluated, yielding
$\langle E^{{\dagger}}(\omega_{b}^{\prime})E^{{\dagger}}(\omega_{a}^{\prime})E(\omega_{b})E(\omega_{b})\rangle_{\rho_{\text{photon}}}$
$=F^{*}(\omega_{b}^{\prime},\omega_{a}^{\prime})F(\omega_{a},\omega_{b})$ (6)
We defined, $F(\omega_{a},\omega_{b})=\alpha A_{p}(\omega_{p})\Big\{\mathrm{sinc}[\phi(\omega_{a},\omega_{b})]$, where $\phi(\omega_{a},\omega_{b})=(\omega_{a}-\omega_{1})\tilde{T}_{1}/2+(\omega_{b}-\omega_{1})\tilde{T}_{2}/2$ and $\alpha$ denotes the parameter governing the down-conversion efficiency. The entanglement time parameter $\tilde{T}_{\mathrm{ent}}=\tilde{T}_{2}-\tilde{T}_{1}$ establishes an upper bound for the intervals between two photon-exciton interactions and depends on the time delays acquired by the signal and idler photons during propagation within the nonlinear crystal. It characterizes the non-classical correlation in the photonic source. The classical transform-limited pulse driving the SPDC crystal is defined as follows: $A_{p}(\omega_{a},\omega_{b})=E_{0}\sqrt{\pi/\Gamma_{p,0}}\exp{[-(\omega-\omega_{a}-\omega_{b})^{2}/4\Gamma_{p,0}]}$, where $\Gamma_{p,0}$ is the temporal width.